Fermi National Laboratory
 

Accelerator Update for November 8, 2006 - November 10, 2006

 

Wednesday November 8

The day shift began with Operations monitoring a Recycler stash (234.5E10), with the Antiproton Source (Pbar) stacking (38.8E10), with Linac, Booster, and the Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for a store, and with MiniBooNE, and NuMI taking beam.

At 8:34 AM, the MI power supply loop beam permit tripped off. Because of this trip, the Recycler lost 150E10 of its stash. Experts investigated and discovered a failed MI-10 power supply door switch.

Operations put the TeV into a dry squeeze at 9:01 AM.

At 10:02 AM, MI resumed ramping, but Pbar's collection lens (lithium lens) had tripped on a LCW conductivity indication. Operations contacted an expert.

NuMI resumed taking beam at 10:07 AM.

Operators cleared the collection lens low conductivity trip. Pbar resumed stacking at 10:15 AM.

At 10:45 AM, a TeV kicker (PAK3) prefired. Operators investigated.

At 11:34 AM, Operations reported that the Pbar collections lens was still tripping. Operations contacted an expert.

Pbar experts reported at 12:32 PM that they found a loose connection on the conductivity meter. Stacking resumed at 12:42 PM.

At 1:42 PM, a Linac RF station (KRF4) tripped off on a cavity LCW flow indication. The cavity temperature returned to normal. Beam resumed at 2:21 PM.

Experts reported that PAK3 was running. Operations put the TeV through a dry squeeze at 2:21 PM.

Operations put the TeV into shot setup at 2:47 PM.

At 4:09 PM, the Pbar lithium lens tripped off again. Operations contacted an expert.

Operations established store 5055 around 4:54 PM with an initial luminosity of 93.07E30.

The Pbar expert reported at 5:06 PM that the lithium lens was up and running.

Pbar resumed stacking at 5:26 PM.

The Recycler completed stashing antiprotons at 6:41 PM.

The Recycler completed stashing at 9:29 PM.

Thursday November 9

The midnight shift began with Operations monitoring store 5055 (41.05E10) and a stash, with Pbar stacking (46.4E10), with Linac, Booster, and MI operating normally, with CDF and D0 using the colliding beams, and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash around 12:24 AM.

The Recycler completed a stash at 3:30 AM.

The I- Source tripped off at 3:47 AM. It reset okay.

At 4:07 AM, a Booster RF station (BRF19) tripped off for the first of three times in twenty-three minutes. After the last trip, operators bypassed the station.

The Recycler completed a stash at 6:25 AM.

At 8:10 AM, store 5055 quenched at TeV sector D1 and F4. Operators investigated and thought that a B11 separator spark had caused the quench.

Operations had CDF and D0 ready for an access by 9:08 AM.

Operations put the TeV into ramp mode at 10:55 AM.

MiniBooNE resumed taking beam at 11:14 AM.

The Recycler completed a stash at 12:07 PM.

Operations put the TeV into shot setup at 12:56 PM.

Operations established store 5057 at 2:49 PM with an initial luminosity of 182.25E30.

Operations prepared the Recycler to receive antiprotons from Pbar at 4:57 PM. The Accumulator RF didn't look normal for the first transfer and the second transfer didn't work at all.

Operations contacted an expert about the stashing problems. Pbar resumed stacking at 5:38 PM.

The Recycler completed a stash at 6:34 PM.

At 6:36 PM, a Booster power supply (GMPS) suffered a transductor fault, which caused a glitch with the low-level RF.

At 6:43 PM, Booster developed a problem with its VB foil and VIsep valves. The Booster vacuum looked fine, but the foil gauge wasn't. This problem continued and Operations called in an expert.

At 8:21 PM, the Booster expert reported that the foil's power supply shorted out. He swapped out the power supply.

Booster resumed sending beam to MiniBooNE 7:27 PM.

At 7:30 PM, the NuMI beam permit tripped on a gas flow indication. Operations contacted experts, followed their instructions, and had beam to NuMI within forty-five minutes.

At 8:49 PM, a Linac Klystron RF station (KRF1) tripped off for the first of six times in eighty minutes on a coil comparator fault. It reset each time, but at 10:11 PM operators contacted an expert for help.

The Recycler completed a stash at 10:41 PM.

Friday November 10

The midnight shift began with Operations monitoring store 5057 (69.93E30) and a stash (76.9E10), with Pbar stacking (35.7E10), with Linac having problems with KRF1, with Booster and MI operating normally, with CDF and D0 using the colliding beams, and with NuMI and MiniBooNE taking beam.

At 12:21 AM, the Linac expert reported a ground fault indication, but it was due to loose connections.

At 1:47 AM, a Klystron Debuncher tripped and wouldn't reset.

The Recycler completed a stash at 2:02 AM. Unfortunately, Pbar couldn't resume stacking due to the Klystron Debuncher problem. Operations called in a Linac expert.

At 3:21 AM, a MI RF station (MIFR5) tripped off on a RF fault and a cavity LCW leak indication. Operators bypassed the station.

At 3:25 AM, the TeV tune began to wander. CDF and D0 turned off their detectors.

At 3:26 AM, Operations discovered that a TeV quadrupole (QF32) had tripped off earlier.

At 3:34 AM, the TeV suffered a quench as operators tried to reset and turn on QF32.

Also at 3:34 AM, the Klystron Debuncher tripped off again. The expert, who was still here, went back to work and had it back on line by 4:05 AM.

At 6 AM, Operations reported a temperature problem with the MI-20 LCW system. An LCW expert investigated and reported that the its pond water strainer was clogged. FESS personnel checked the strainer.

Operations put the TeV through a dry squeeze at 7:21 AM.

Operations put the TeV into shot setup at 7:50 AM.

FESS reported at 7:54 AM that the MI-20 pond strainer had been cleaned.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/10/2006    email Fermilab